how many grams of Cl are in 265g CaCl2 ?
1
Expert's answer
2019-09-23T08:19:56-0400

At the first, we need to find the mass fraction of "{{\text{Cl}}}" in "{{\text{CaC}}{{\text{l}}_{\text{2}}}}". To do so we need to find the molar mass of "{{\text{CaC}}{{\text{l}}_{\text{2}}}}" - it can be calculates as the sum of atomic weights taking into account the number of atoms of each type.


"{M_{{\text{CaC}}{{\text{l}}_{\text{2}}}}} = {M_{{\text{Ca}}}} + 2{M_{{\text{Cl}}}} = 40.08[\frac{{\text{g}}}{{{\text{mol}}}}] + 2 \cdot 35.45[\frac{{\text{g}}}{{{\text{mol}}}}] = 110.98[\frac{{\text{g}}}{{{\text{mol}}}}]"

(atomic weights of atoms "{{\text{Ca}}}" and "{{\text{Cl}}}" can be found in periodic table and respectively equal to "40.08[\frac{{\text{g}}}{{{\text{mol}}}}]" and "35.45[\frac{{\text{g}}}{{{\text{mol}}}}]")

Then we can find the mass fraction by dividing the sum of atomic weights of all "{{\text{Cl}}}" atoms by molecular mass of "{{\text{CaC}}{{\text{l}}_{\text{2}}}}"


"\omega = \frac{{2{M_{{\text{Cl}}}}}}{{{M_{{\text{CaC}}{{\text{l}}_{\text{2}}}}}}} = \frac{{2 \cdot 35.45[\frac{{\text{g}}}{{{\text{mol}}}}]}}{{110.98[\frac{{\text{g}}}{{{\text{mol}}}}]}} = 0.64"


And let's find the mass of "{{\text{Cl}}}" by multiplying the mass fraction by mass of "{{\text{CaC}}{{\text{l}}_{\text{2}}}}"


"{m_{{\text{Cl}}}} = \omega {m_{{\text{CaC}}{{\text{l}}_{\text{2}}}}} = 0.64 \cdot 265[{\text{g}}] = 169.6[{\text{g}}]"